[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Oct 18, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Wed Oct 18 14:20:23 UTC 2017


Here is today's summary of NHC computer operations:
--- Mello

----------

TAFB:
1. Doug indicated NCO will roll back the NAWIPS red hat 7 boxes back to red  
hat 6 that should fix the scatterometer display problems.

2. TAFB situation awareness monitor W wall locked up. Recycled, but  
previous login's failed. MN/MH. Marc L will follow up

3. Mello fixed the Buoy 41046 location in D2D.
HSU:
1. There was a Goes-West Product Outage/Anomaly last night before 0127Z.  
Mike Brennan reported that the ESPC was informed by the GOES controller and  
that the wide band issue was resolved.
Other items:

NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171018/c254f416/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list